# Planets and planetesimals at cosmic dawn: Vortices as planetary nurseries

### v1

@@ 4 Formation of the first planets
Figure 3: Final planetary masses for the same simulation that is presented in Fig. 1, for different assumptions of the embryos residence time within the vortex. The red dotted line outlines the region where planetesimals form in case VC.
In this section, we address whether the planetesimals that form are likely to act as seeds for planet formation, or if yet higher metallicities are required for planetary objects to form. To save time and computational resources, we use a coarser space-time grid in this part of the study. Fig. 3 shows the final planetary masses for the same simulation that is presented in Fig. 1, for different assumptions of the embryos residence time within the vortex.
@@
The bottom plot of Fig. 3 shows the final planetary masses when the embryos reside inside the vortex for 100 orbital periods. In this scenario, the embryo accretes all of the trapped dust and experiences an additional 99 orbital periods of vortex-assisted pebble accretion. This results in a $\sim 15-30\%$ increase in the final planetary masses compared to when the residence time was one orbital period. These results show that the accretion of trapped dust is the most important contributor to planet formation inside vortices in low-metallicity disks, with subsequent pebble accretion playing a lesser role.
Figure 4: The plots show the final planetary masses for all simulations in the parameter study. In these simulations it was assumed that the embryos reside ten orbital periods within the vortex, after which they continue to accrete at the normal accretion rate.
Fig. 4 displays the final planetary masses for all simulations in the parameter study, assuming a residence time of ten orbital periods within the vortex. Generally, larger $Z$ and lower $\alpha_{\rm t}$ results in larger planetary masses. The largest planets that are able to form in our study have a mass of $\sim 0.7\,\textrm{M}_{\oplus}$, and form when $Z=1.6\times 10^{-3}$ and $\alpha_{\rm t}=10^{-5}$. When $Z=4\times 10^{-4}$ and $\alpha_{\rm t}=10^{-5}$, we obtain planets with masses just below that of Mercury. We note that these results are heavily dependent on the calculation of the trapped dust mass within the vortex (Eq. 7), and the assumption that all trapped dust is accreted by the planet.
Figure 5: The plots display the final planetary masses as a function of the embryos’ formation locations for a selection of simulations shown in Fig. 4. The color of the scatter points represents the formation time of the embryos.
Fig. 5 shows the final planetary masses as a function of semimajor axis for four different simulations. The planetary masses increase with distance from the star due to the strong dependence of trapped dust mass on semimajor axis ($M_{\rm trapped}\propto H^{3}$). If growth were dominated by standard pebble accretion outside the vortex, this trend would be reversed, as the pebble accretion rate decreases with semimajor axis. On average, planets in our simulations attain lower masses the later they form. This is primarily because the background dust-to-gas ratio peaks early in the disk’s evolution and decreases over time. Fig. 5 highlights that the timing and location of vortex formation within the disk significantly impact the masses of the planets forming inside it.
@@ 5.1 Comparison with previous work
In this study, we investigate whether vortices can trigger the formation of the first planets and planetesimals, and at what metallicity this could have occurred. Although the exact metallicity threshold varies with the assumed model and model parameters, most of our simulations suggest that planetesimals and Mercury-mass planets can form inside vortices when $Z\gtrsim 0.04Z_{\odot}$, provided a low level of turbulence. Previous studies on the critical metallicity threshold for planet formation include Johansen et al. 2009, Johnson & Li (2012), and Carrera et al. (2015).
@@ 5.2.1 Mono-disperse VS poly-disperse size distribution
Regarding the SI criteria used in this work, both are derived from simulations using a mono-disperse size distribution. Linear studies of the SI with a poly-disperse size distribution show that the SI is sometimes suppressed (Krapp et al., 2019); however, Zhu & Yang (2021) and Yang & Zhu (2021) found that it can still operate at high dust-to-gas ratios or $\mathrm{St}$. Given the large grain sizes and dust-to-gas ratios expected at the vortex center, it is reasonable to assume that the SI can still operate. How the SI criteria change from the mono-disperse to the poly-disperse case remains an active area of research.
@@ Acknowledgements
LE acknowledges the support from NASA via the Emerging Worlds program (#80NSSC25K7117), as well as funding by the Institute for Advanced Computational Science Postdoctoral Fellowship. WL acknowledges support from the NASA Theoretical and Computational Astrophysical Networks (TCAN) via grant #80NSSC21K0497. WL is further supported by grants #80NSSC22K1419 from the NASA Emerging Worlds program and AST-2007422 from NSF. DC and WL acknowledge support from NASA under Emerging Worlds through grant 80NSSC25K7414. B.B. acknowledges support from NSF grant AST-2009679 and 2407877. B.B. and S.M. acknowledge support from NASA grant No. 80NSSC20K0500. B.B. is grateful for generous support from the David and Lucile Packard Foundation and the Alfred P. Sloan Foundation.
@@
The data underlying this article will be shared on reasonable request to the corresponding author.
@@ Appendix A Calculation of residence time inside the vortex
The vortex can be described by an ellipse with semi-minor axis $a$ and semimajor axis $b=a\chi$, where $\chi=4$ and $a=2H/3$ for the Kida solution. Let the vortex center have coordinates $(x,y)=(0,0)$, where x points in the direction of the semi-minor axis (the radial direction) and y points in the direction of the semimajor axis (the direction along the orbit). The equation for an ellipse is:
@@ Appendix B Increasing the vortex lifetime
Figure 7: The plots show how the planetesimal formation regions in case VCG changes when different vortex lifetimes are assumed. These results are obtained using SI criteria SIa.
@@ Appendix C Simulations with a solar mass star and varying disk-to-star mass ratios
Figure 8: The plots show the locations for planetesimal formation inside vortices when considering a stellar mass of $1\,\textrm{M}_{\odot}$ and a disk-to-star mass ratio of 10% (left plot) and 1% (right plot).
Figure 9: The plots show the final planetary masses for the same planetesimal formation regions shown in Fig 8, when assuming a residence time of 10 orbital periods within the vortex.
@@ Appendix D Dust growth in vortices
We adopt the model of Carrera et al. (in prep), who found a powerful feedback loop between vortex concentration and dust coagulation. We refer to that paper for the details of the model. Here we simply summarize the key ideas and describe the final algorithm. Vortices have several properties that make them prime candidates for planet formation:
@@
  • Turbulence dampening: Vortices are already known to have lower turbulence than their surroundings, independent of dust mass loading (Lesur & Papaloizou, 2010b; Lyra & Klahr, 2011b). The key insight of Carrera et al. (in prep) is that the amount of turbulence dampening is a dynamic quantity, that responds to grain growth occurring within the vortex.
@@
  • Turbulence dampening: Carrera et al. (in prep) derived the turbulent gas velocity, $V_{\mathrm{gas}}$, in a gas-dust mixture where both $\mathrm{St}$ and the midplane dust-to-gas ratio $\epsilon$ are non-negligible:
@@
  • Dust sedimentation: Lastly, mass loading also enhances dust sedimentation. This was seen already in the numerical simulations of Lim et al. (2024b). Again, we use the expression derived by Carrera et al. (in prep):
@@
Lastly, Carrera et al. (in prep) derived the vortex concentration timescale $t_{Z}=1/(\mathrm{St}\Omega)$, but they noted that numerical simulations suggest a slightly higher $t_{Z}$. Therefore, we set
@@ Appendix E Streaming Instability Criterion
The SI criterion from Lim et al. (2024a) (SIb) is expressed in terms of the dust-go-gas surface density:
@@
To use this criterion, we must convert it into the equivalent midplane dust-to-gas density $\epsilon_{\rm crit}$. We can write the midplane density as $\epsilon\approx Z(H/H_{\mathrm{dust}})$, where $H_{\mathrm{dust}}$ is the scale height of the dust layer. Lim et al. (2024b) found that
@@
where $\Pi$ is a parameter that quantifies the strength of the pressure gradient and $\tilde{H}=H/\sqrt{1+\epsilon}$ is the “effective” scale height of the gas-dust mixture. The $(\Pi/5)^{2}$ term is an estimate of the self-generated turbulence due to the SI. The simulations of Lim et al. (2024a) had $\Pi=0.05$ and no forced turbulence. Therefore, their simulations had
@@
At this point we deviate slightly from Lim et al. (2024b). Their expression for $\tilde{H}$ was based on approximating the gas-dust mixture as a colloid ($\mathrm{St}\ll 1$), but we are interested in the case where $\mathrm{St}$ is not small. Therefore, we use the formula from Carrera et al. (in prep), who derived $\tilde{H}$ without the colloid approximation,
@@
Notice that, in the limit as $\mathrm{St}\rightarrow 0$, this expression becomes identical to the one in Lim et al. (2024b). Considering the small $\mathrm{St}$ values in their simulations, both expressions are equally consistent with their dataset. Altogether, the midplane dust-to-gas ratio is

### v2

@@ 4 Formation of the first planets
Figure 3: Final masses of planets forming inside a vortex as a function of the formation time and semimajor axis of the vortex, for the same simulation presented in Fig. 1. The three panels employ different assumptions for the embryos’ residence time within the vortex. The results shown are for case VCG, with the red dotted line outlining the region where planetesimals form in case VC.
In this section, we address whether the embryos that form are likely to act as seeds for planet formation, or if even higher metallicities are required for planetary objects to form. To save time and computational resources, we use a coarser space-time grid in this part of the study. For each point on the grid where vortex formation would result in planetesimal formation, we calculate the embryo mass (taken to be 10x the typical planetesimal mass) and simulate growth via pebble accretion onto this embryo. Fig. 3 shows the resulting planetary masses at the end of the disk lifetime for the same simulation presented in Fig. 1. The different panels employ different assumptions for the embryos’ residence time within the vortex. The results should be interpreted as follows: if a vortex were to form at $(r=r_{j},t=t_{i})$, the embryo that forms inside would grow to a mass $M_{\rm p}(j,i)$ by the end of the disk lifetime.
@@
[…]
Figure 4: The plots show the final masses of planets forming inside a vortex as a function of the formation time and semimajor axis of the vortex, for all simulations in the parameter study. In these simulations it was assumed that the embryos reside ten orbital periods within the vortex, after which they continue to accrete at the normal accretion rate.
Fig. 4 shows the planetary masses at the end of the disk lifetime for all simulations in the parameter study, assuming a residence time of ten orbital periods within the vortex. Generally, vortices that form in disks with higher $Z$ and lower $\alpha_{\rm t}$ produce planets with larger masses. The largest planets formed in our study have masses of $\sim 0.7\,\textrm{M}_{\oplus}$ and occur when $Z=1.6\times 10^{-3}$ and $\alpha_{\rm t}=10^{-5}$. When $Z=4\times 10^{-4}$ and $\alpha_{\rm t}=10^{-5}$, we obtain planets with masses just below that of Mercury. We note that these results are heavily dependent on the calculation of the trapped dust mass within the vortex (Eq. 7) and the assumption that all trapped dust is accreted by the planet.
Figure 5: The plots display the final planetary masses as a function of the vortex formation location for a selection of the simulations shown in Fig. 4. The color of the scatter points represents the formation time of the vortex.
Fig. 5 shows the masses of planets that are able to form as a function of the vortex formation location, for four different simulations. The planetary masses increase with the vortex’s formation distance from the star due to the strong dependence of trapped dust mass on semimajor axis ($M_{\rm trapped}\propto H^{3}$). If growth were dominated by standard pebble accretion outside the vortex, this trend would be reversed, as the pebble accretion rate decreases with semimajor axis. On average, planetary masses decrease with increasing vortex formation time. This is primarily because the background dust density peaks early in the disk’s evolution and decreases over time, resulting in vortices that form later having a smaller dust budget. Fig. 5 highlights that the timing and location of vortex formation within the disk significantly impact the masses of the planets forming inside it.
@@ 5.1 Comparison with previous work
In this study, we investigate whether vortices can trigger the formation of the first planets and planetesimals, and at what metallicity this could have occurred. Although the exact metallicity threshold varies with the assumed model and model parameters, most of our simulations suggest that planetesimals and Mercury-mass planets can form inside vortices when $Z\gtrsim 0.04Z_{\odot}$, provided a low level of turbulence. Previous studies on the critical metallicity threshold for planet formation include Johansen et al. 2009, Johnson & Li (2012), Carrera et al. (2015) and Andama et al. (2024).
@@ 5.2.1 Mono-disperse VS poly-disperse size distribution
Regarding the SI criteria used in this work, both are derived from simulations using a mono-disperse size distribution. Linear studies of the SI with a poly-disperse size distribution show that the SI is sometimes suppressed (Krapp et al., 2019); however, Zhu & Yang (2021) and Yang & Zhu (2021) found that it can still operate at high dust-to-gas ratios or $\mathrm{St}$. More recently, Ho et al. (2024) presented simulations of the polydisperse SI with dust coagulation taken into account. They find that the interplay between dust coagulation and the SI can result in planetesimal formation even for relatively small initial dust sizes. Given the large grain sizes and dust-to-gas ratios expected at the vortex center, it is reasonable to assume that the SI can still operate. How the SI criteria change from the mono-disperse to the poly-disperse case remains an active area of research.
@@ Acknowledgements
The authors wish to thank the anonymous referee for providing helpful comments that lead to an improved manuscript. LE acknowledges the support from NASA via the Emerging Worlds program (#80NSSC25K7117), as well as funding by the Institute for Advanced Computational Science Postdoctoral Fellowship. WL acknowledges support from the NASA Theoretical and Computational Astrophysical Networks (TCAN) via grant #80NSSC21K0497. WL is further supported by grants #80NSSC22K1419 from the NASA Emerging Worlds program and AST-2007422 from NSF. DC and WL acknowledge support from NASA under Emerging Worlds through grant 80NSSC25K7414 and Exoplanets Research Program through grant 80NSSC24K0959. B.B. acknowledges support from NSF grant AST-2009679 and 2407877. B.B. and S.M. acknowledge support from NASA grant No. 80NSSC20K0500. B.B. is grateful for generous support from the David and Lucile Packard Foundation and the Alfred P. Sloan Foundation.
@@
All simulations and results can be reproduced using our open-source code firstPlanets, available on GitHub: https://github.com/astrolinn/firstPlanets. Furthermore, the data underlying this article will be shared on reasonable request to the corresponding author.
@@ Appendix A Calculation of residence time inside the vortex
[…]
@@ Appendix B Increasing the vortex lifetime
Figure 7: The plots show how the region where vortex formation results in planetesimal formation changes when different vortex lifetimes are assumed. These results are obtained using SI criteria SI24.
@@ Appendix C Simulations with a solar mass star and varying disk-to-star mass ratios
Figure 8: The plots show where on the space-time grid vortex formation results in planetesimal formation, when a stellar mass of $1\,\textrm{M}_{\odot}$ and a disk-to-star mass ratio of 10% (left plot) and 1% (right plot) are adopted.
Figure 9: The plots show the final planetary masses for the embryos formed in Fig 8, when assuming a residence time of 10 orbital periods within the vortex.
@@ Appendix D Dust growth in vortices
We adopt the model of Carrera et al. (2025), who found a powerful feedback loop between vortex concentration and dust coagulation. We refer to that paper for the details of the model. Here we simply summarize the key ideas and describe the final algorithm. Vortices have several properties that make them prime candidates for planet formation:
@@
  • Turbulence dampening: Vortices are already known to have lower turbulence than their surroundings, independent of dust mass loading (Lesur & Papaloizou, 2010b; Lyra & Klahr, 2011b). The key insight of Carrera et al. (2025)) is that the amount of turbulence dampening is a dynamic quantity, that responds to grain growth occurring within the vortex.
@@
  • Turbulence dampening: Carrera et al. (2025) derived the turbulent gas velocity, $V_{\mathrm{gas}}$, in a gas-dust mixture where both $\mathrm{St}$ and the midplane dust-to-gas ratio $\epsilon$ are non-negligible:
@@
  • Dust sedimentation: Lastly, mass loading also enhances dust sedimentation. This was seen already in the numerical simulations of Lim et al. (2024). Again, we use the expression derived by Carrera et al. (2025):
@@
Lastly, Carrera et al. (2025) derived the vortex concentration timescale $t_{Z}=1/(\mathrm{St}\Omega)$, but they noted that numerical simulations suggest a slightly higher $t_{Z}$. Therefore, we set
@@ Appendix E Streaming Instability Criterion
The SI criterion from Lim et al. (2025) (SI25) is expressed in terms of the dust-go-gas surface density:
@@
To use this criterion, we must convert it into the equivalent midplane dust-to-gas density $\epsilon_{\rm crit}$. We can write the midplane density as $\epsilon\approx Z(H/H_{\mathrm{dust}})$, where $H_{\mathrm{dust}}$ is the scale height of the dust layer. Lim et al. (2024) found that
@@
where $\Pi$ is a parameter that quantifies the strength of the pressure gradient and $\tilde{H}=H/\sqrt{1+\epsilon}$ is the “effective” scale height of the gas-dust mixture. The $(\Pi/5)^{2}$ term is an estimate of the self-generated turbulence due to the SI. The simulations of Lim et al. (2025) had $\Pi=0.05$ and no forced turbulence. Therefore, their simulations had
@@
At this point we deviate slightly from Lim et al. (2024). Their expression for $\tilde{H}$ was based on approximating the gas-dust mixture as a colloid ($\mathrm{St}\ll 1$), but we are interested in the case where $\mathrm{St}$ is not small. Therefore, we use the formula from Carrera et al. (2025), who derived $\tilde{H}$ without the colloid approximation,
@@
Notice that, in the limit as $\mathrm{St}\rightarrow 0$, this expression becomes identical to the one in Lim et al. (2024). Considering the small $\mathrm{St}$ values in their simulations, both expressions are equally consistent with their dataset. Altogether, the midplane dust-to-gas ratio is